# Can the Waymo Open Motion Dataset Support Realistic Behavioral Modeling? A Validation Study with Naturalistic Trajectories

### v1

@@ 5.1 Identification of Lane-changing cases
Third, we assigned lead and lag vehicles to all vehicles at each time step to support the gap acceptance analysis described in the next section. In urban scenarios, where center lanes often split into multiple successors (e.g., a shared right-turn/straight lane), multiple leader candidates may exist. We selected the leader with the shortest longitudinal distance along the assigned center line. Figure 10 shows an example trajectory of a lane-changing vehicle along with its lead and lag vehicles.
@@
The observed mean difference was found to be $\bar{T}^{\text{LC}}_{\text{obs}}=0.1803$. The permutation test yielded a $p$-value of $p=0.0086$, indicating that the cross-set distances are statistically larger than the internal variability of WOMD. Figure 11 visualizes the null distribution of mean differences and the observed test statistic. These results suggest that the lane-changing dynamics observed in PHX do not lie within the behavioral envelope of WOMD, despite accounting for distance and velocity error in the PHX measurements. Thus, caution is warranted when using WOMD to model lane-changing behavior.
@@ 6 Conclusion
This study critically examined the validity of the usage of WOMD in driving behavior analysis by comparing its trajectory data to an independently collected naturalistic dataset from Level 4 AV operations in Phoenix, AZ (PHX). We focused on three scenarios: intersection discharging, car-following, and lane-changing. Using manual annotation from high-resolution aerial video in the PHX dataset, we established a reliable reference for ground-truth headways at intersections. For car-following and lane-changing, we applied a SIMEX correction to account for observational noise in the PHX data and used DTW distance to measure behavioral differences.
Across all scenarios, we found a significant difference between the behaviors observed in WOMD and those captured in PHX. WOMD consistently underrepresents short headways and sharp decelerations. These differences exist even after correcting for measurement error in PHX, suggesting that the differences stem not from data quality but from underlying behavioral characteristics. Permutation tests confirm that these cross-dataset differences are statistically significant and lie outside the internal variability of WOMD.

### v2

@@ 5.1 Identification of Lane-changing cases
Third, we assigned lead and lag vehicles to all vehicles at each time step to support the gap acceptance analysis described in the next section. In urban scenarios, where center lanes often split into multiple successors, for example, at a shared right-turn/straight lane, multiple leader candidates may exist. We selected the leader with the shortest longitudinal distance along the assigned center line. Figure 10 shows an example trajectory of a lane-changing vehicle along with its lead and lag vehicles. The plot shows a $6$-second window centered at the frame when the lane-changing vehicle crosses the lane marking ($t=0$). Longitudinal position is measured along the lane centerline, and lateral position represents cross-lane displacement relative to the initial lane.
@@
We applied the same outlier filter as in the car-following analysis: lane changes whose six-second window contained longitudinal accelerations larger than $10~\mathrm{m/s^{2}}$ in magnitude or spacing jumps greater than $30~\mathrm{m}$ over $0.1$ s were excluded. The final sample consists of 13 PHX lane-change episodes and 200 comparator episodes from WOMD. In the PHX dataset, the ego vehicle in all analyzed lane-change events is a commercial robotaxi, while all surrounding vehicles are human-driven. During our observation period, we did not encounter any lane changes in which the lead or lag vehicle in the target lane was also an automated vehicle. Consequently, our lane-changing analysis reflects mixed traffic with a single AV ego rather than AV–AV interactions, which remain rare in current deployments.
@@
The observed mean difference was found to be $\bar{T}^{\text{LC}}_{\text{obs}}=0.1803$. The permutation test yielded a $p$-value of $p=0.0086$, indicating that the cross-set distances are statistically larger than the internal variability of WOMD. The main message is that several everyday lane changes observed in PHX are more dissimilar to any WOMD lane change than typical WOMD lane changes are to one another. We emphasize that this conclusion is based on 13 lane-changing episodes from PHX and 200 comparator episodes from WOMD; while the permutation test p-value indicates a statistically detectable difference at this sample size, the effect size and qualitative shape of the distance distributions are more informative than the exact p-value. Figure 11 visualizes the null distribution of mean differences and the observed test statistic. These results suggest that the lane-changing dynamics observed in PHX do not lie within the behavioral envelope of WOMD, despite accounting for distance and velocity error in the PHX measurements. Thus, caution is warranted when using WOMD to model lane-changing behavior.
@@ 6 Conclusion
This study introduced a robust methodology for validating trajectory datasets using an independent naturalistic ground-truth dataset with quantified error distribution. By combining manual, frame-anchored ground truth for intersection discharge headways and SIMEX-based correction and distributional tests for car-following and lane-changing behaviors, the framework enables validity claims that do not depend on any post-processing of the public dataset. This independent validation perspective complements earlier trajectory quality studies that assessed datasets such as NGSIM primarily through internal diagnostics and filtered-trajectory references.
Particularly, this study critically examined the extent to which WOMD can be used in driving behavior analysis by comparing its trajectory data to an independently collected dataset of commercial robotaxi operations in Phoenix, AZ (PHX). We focused on three scenarios: intersection discharging, decelerating-to-stop, and lane-changing. Because the two datasets differ in collection period, operational design domain, and sensing modality, any discrepancies we observe necessarily conflate differences in control policy, roadway environment, and post-processing pipelines. Our analysis, therefore, does not aim to isolate a single causal factor, but instead asks a more practical question: do the behaviors observed in recent, naturalistic Level 4 operations fall inside the behavioral envelope spanned by WOMD trajectories?
Across all scenarios, we found a significant difference between the behaviors observed in WOMD and those captured in PHX. WOMD consistently underrepresents short headways and sharp decelerations. Because the PHX discharge headways are anchored in frame-by-frame manual annotations of vehicles crossing the stop line, with uncertainty at the level of a single video frame, these discrepancies cannot be explained away as artefacts of aerial measurement error; they instead point to genuine differences between the naturalistic acceleration behavior we observe in recent robotaxi operations and the start-up dynamics represented in WOMD. These differences exist even after correcting for measurement error in PHX, suggesting that the differences stem not from data quality but from underlying behavioral characteristics. Permutation tests confirm that these cross-dataset differences are statistically significant and lie outside the internal variability of WOMD.
@@
More broadly, our study illustrates both the value and the limitations of cross-source validation for AV trajectory datasets. Because PHX and WOMD differ in operational design domain, geography, time period, and sensing stack, we can only conclude that certain patterns observed in recent naturalistic operations, such as short headways and sharp decelerations, are underrepresented in the WOMD dataset. We cannot uniquely attribute the discrepancy to a specific layer of the AV perception stack or to a particular post-processing choice. Future validation efforts will benefit from coordinated data releases that include both raw sensor streams and independent external measurements in overlapping time and space.